OpenStudy (anonymous):

solve each system by substitution. -2x-3y=-9 2x-2y=4

OpenStudy (anonymous):

x = 15 y = 13

OpenStudy (anonymous):

Alright, we have to get y by itself on one equation, let's do 2x-2y=4 -2x -2x Subtract 2x from both sides to get y by itself. -2y=4 - 2x /-2 /-2 /-2 Divide by -2 from both sides to fully get y by itself. now we know that y=-2+x So we can fit that into y's spot on the other equation -2x-3y=-9 -2x-3(-2+x)=-9 Distribute -3, so -3(-2) and (-3)(x) -2x+6-3x=-9 <Now we've got like terms, add them together -5x+6=-9 -6 -6 Subtract 6 from both sides -5x=-15 /-5 /-5 Divide by -5 on both sides x=3

OpenStudy (anonymous):

Then to find out what y=, you just fit x into the problem, so 2x-2y=4 3(3) -2y=4 9-2y=4 +9 +9 -2y=4 /-2 /-2 y=-2

OpenStudy (anonymous):

thanks ess, that took a while to write out, lol.

OpenStudy (anonymous):

lol i can tell.